OpenStudy (anonymous):

What is the equation of a line that passes through (-6, 2) and has a slope of −1/2 ?

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

where m= slope, y1= y coordinate, and x1 = x coordinate

OpenStudy (anonymous):

plug in the information you have been given and solve for y

OpenStudy (anonymous):

Hi :) Slope-intercept of a line: \(\LARGE y=mx+b\) where m=slope, be is the y-intercept. Standard equation of a line: \(\LARGE Ax+By=C\) which form do you need?

OpenStudy (anonymous):

this is the choices ..A) y=x2−2 B) y=−x2−1 C) y=2x+1 D )y=−2x−1

OpenStudy (anonymous):

i would recommend point slope form then

OpenStudy (anonymous):

Use the point slope form \(y-y _{1}=m(x-x _{1}) \) Put this into slope intercept form \(y-2=\frac{1}{2}(x-(-6)) \)

OpenStudy (anonymous):

yes, like @MDoodler from there, solve for y

OpenStudy (anonymous):

I'm lost..

OpenStudy (anonymous):

alright i will try to explain it.

OpenStudy (anonymous):

point slope form uses a given point {in your case (-6, 2)}

OpenStudy (anonymous):

and the slope (in your case -1/2)

OpenStudy (anonymous):

and by using \[y-y _{1}=m(x-x _{1})\] you can find your equation in the form or y=mx + b which is called standard form

OpenStudy (anonymous):

so in your case, -6 = your x1 2 is your y1 and -1/2 is your m

OpenStudy (anonymous):

try putting those values into your equation

OpenStudy (anonymous):

y = mx + b is not standard form TechnoSoul.l That form is slope intercept form.

OpenStudy (anonymous):

you should end up with \[y-2=-1/2(x+6)\]

OpenStudy (anonymous):

sorry, ya. slope-intercept form. my bad

OpenStudy (anonymous):

Its fine I think I get it now

OpenStudy (anonymous):

alright, once you get your answer, you can write it here and i can check that its correct if you want

OpenStudy (anonymous):

so I times -1/2 and 6?

OpenStudy (anonymous):

yes good

OpenStudy (anonymous):

by now, you should have something that looks like \[y-2=-1/2x-3\]

OpenStudy (anonymous):

if you dont have that i can explain what i did

OpenStudy (anonymous):

and then from there add 2 to both sides and you have your equation

OpenStudy (anonymous):

okay i got y=1/2x-1

OpenStudy (anonymous):

wait its y=x/2-1 right?

OpenStudy (anonymous):

yes it is good job :)

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

sure thing